### v1

@@ Abstract
Urban morphology has long been recognized as a factor shaping human mobility, yet comparative and formal classifications of urban form across metropolitan areas remain limited. Building on theoretical principles of urban structure and advances in unsupervised learning, we systematically classified the built environment of nine U.S. metropolitan areas using structural indicators such as density, connectivity, and spatial configuration. The resulting morphological types were linked to mobility patterns through descriptive statistics, marginal effects estimation, and post hoc statistical testing. Here we show that distinct urban forms are systematically associated with different mobility behaviors, such as reticular morphologies being linked to significantly higher public transport use (marginal effect = 0.49) and reduced car dependence (–0.41), while organic forms are associated with increased car usage (0.44), and substantial declines in public transport (–0.47) and active mobility (–0.30). These effects are statistically robust (), highlighting that the spatial configuration of urban areas plays a fundamental role in shaping transportation choices. Our findings extend previous work by offering a reproducible framework for classifying urban form and demonstrate the added value of morphological analysis in comparative urban research. These results suggest that urban form should be treated as a key variable in mobility planning and provide empirical support for incorporating spatial typologies into sustainable urban policy design.
@@
The physical structure of cities exerts a determining influence on the daily lives of their inhabitants, shaping the way they move, access services, and participate in urban dynamics. In this context, urban morphology stands as a crucial factor in the organization of mobility. This morphology not only defines the accessibility and connectivity of a territory but also conditions the modal share of trips and, consequently, equitable access to the city’s work, educational, and social opportunities (Geurs and Wee, 2004).
Historically, urban and transport policies have tended to address mobility from a predominantly functionalist perspective, focused on the operational efficiency of transport systems. However, this approach has often underestimated how urban design can either foster or restrict certain modes of travel. Faced with pressing contemporary challenges such as vehicular congestion, spatial fragmentation, inequity in access, and the environmental impact of transport, understanding the relationship between urban morphology and mobility is crucial. Prior studies have demonstrated that some urban forms promote active travel and public transit, whereas others foster car dependency (Ewing and Cervero, 2010).
In this context, the present study contributes to understanding how the morphological configuration of street networks influences urban mobility patterns. It examines the structural attributes of road systems as captured by topological and spatial indicators (Bamakan et al., 2019; Cardillo et al., 2006), and explores the extent to which variations in network form are associated with differences in modal share across U.S. cities, providing empirical insights that may support more integrated approaches to urban and transport planning. Through a comparative approach, it aims to provide empirical evidence that can guide the design of more integrated urban and transport planning policies, in order to promote more equitable, sustainable, and efficient mobility systems.
@@ 2 Urban Morphology and Spatial Structure
In this study, we adopt a network-based understanding of urban morphology, where the structure and geometry of the street network are treated as a proxy for the underlying form of the city (Porta et al., 2006; Boeing, 2019). We acknowledge that this is only one of many valid ways to study urban form and do not claim conceptual primacy. Rather, we recognize that form is multidimensional, and that our approach captures one of its structural expressions—specifically, the topological and geometric configuration of urban streets as they relate to patterns of movement and accessibility.
@@ 2.1 Characterization of Urban Morphology: Fundamental Metrics and Dimensions
From this perspective, centrality metrics provide an additional layer of analysis that enables the examination of how accessibility and connectivity are distributed across different street fabrics. When interpreted through a functional lens, these measures help identify not only local connection patterns but also global articulation dynamics within the network. Degree centrality (Freeman, 1978), by estimating the number of direct connections each node has relative to the total, offers a first approximation of its level of integration; building on this, closeness centrality (Musiał et al., 2009; Wasserman and Faust, 1994) broadens the scope by considering the average distances between nodes, allowing for the identification of strategic points with greater potential access to the overall system. Complementarily, betweenness centrality as elaborated by Freeman (1978), introduces a critical dimension by revealing which nodes tend to lie along the shortest paths between other pairs, highlighting their role as articulators of general flow. A similar logic can be applied to the network’s segments through edge betweenness centrality (Lu and Zhang, 2013), which helps identify links with a high structural weight in overall connectivity. Although these metrics were not part of the empirical analysis developed in this study, they represent a complementary analytical dimension that can enrich the study of spatial configurations and the structural understanding of urban networks.
@@ 2.2.4 Hybrid
3 Classification and Treatment of Urban Mobility
In order to analyze the relationship between urban morphology and mobility patterns, data on street space usage are collected and classified based on information provided by the U.S. Census Bureau (2020). This information is organized following the ABC of Mobility framework Prieto-Curiel and Ospina (2024), which classifies travel modes into three categories based on their function and means of transport. Active mobility includes non-motorized forms such as walking and cycling, public mobility comprises trips made using shared or collective transport systems, and private mobility refers to the use of individual motorized vehicles. This classification serves to structure the analysis of modal distribution in a clear and comparable way. Based on this classification, mobility patterns are examined from a dual perspective. On the one hand, the disaggregated approach makes it possible to assess the specific contribution of each transport mode and its potential correlation with particular morphological attributes. On the other hand, the aggregated analysis provides a general overview of the modal composition of each urban unit, allowing for systematic comparisons across different spatial contexts. This methodological approach not only enables the characterization of transport mode distribution but also facilitates the exploration of potential associations between the built environment and everyday mobility practices.
@@ 4 Methodology for Theoretical Classification and Clustering of Urban Patterns
To conduct the morphological analysis of the street networks in the selected cities, the urban territory was divided into administrative units defined by the Census Bureau, known as census tracts. This territorial segmentation enables a precise and consistent characterization of urban morphology, facilitating the integration of topological and spatial properties with complementary data such as mobility indicators available at the tract level. This approach enables systematic comparisons across urban areas while preserving the granularity needed to capture internal heterogeneity.
Urban morphology are analyzed based on the variables previously defined and summarized in Table LABEL:tab:urban_metrics, which group key spatial and topological characteristics relevant to describing street networks. Building on this foundation, a systematic and replicable classification method is applied to quantify the degree of correspondence between each urban unit and the theoretical profiles of urban patterns established in the literature (Section 2.2). This method employs a Multi-Attribute Decision Making (MADM) (Triantaphyllou, 2000) framework that assigns weighted scores and penalties according to how closely the observed morphological attributes align with the characteristic values of each pattern. Within this framework, three types of intervals are identified. The optimal interval includes values that are highly characteristic and representative of a specific pattern, the moderate interval comprises values compatible with the pattern but less distinctive or potentially overlapping with others, and the critical or penalizing interval encompasses values that significantly contradict the defining features of the pattern. For example, a very high proportion of dead-end streets is considered critical when evaluating the Gridiron pattern, triggering a penalty within the scoring scheme. To quantify the correspondence between an urban area and each profile, a weighted scoring mechanism with penalties is applied. Let denote the value of dimension , and the weight assigned based on its relevance to the pattern. The partial score for each dimension is determined by the interval in which falls.
@@ 4.2 Cluster-Based Pattern and Subpattern Classification
Although Table LABEL:tab:urban_metrics defined conceptual groups to organize morphological properties, the PCA analysis shows that the empirical expression of these categories may vary depending on urban scale and internal heterogeneity. Nevertheless, the extracted dimensions tend to preserve the structural logic of the theoretical classification, suggesting a robust correspondence between conceptual patterns and actual urban configurations, without implying a rigid segmentation. Based on these dimensions, clustering techniques were applied to identify predominant morphological patterns, selecting the optimal number of clusters using metrics such as the silhouette score. Although the resulting groupings reflect configurations consistent with the theoretical framework, they should be understood as flexible structures in complex urban environments such as those with large territorial extensions or multilevel developments. Peripheral subgroups may emerge that, despite deviating from the cluster centroid, retain fundamental structural properties that justify their classification.
@@ 5.1 Morphological Pattern Characterization in Selected Cities
Figure 2 presents selected examples of morphological categories and sub-patterns in cities like Sant Lake City, Boston, and Chandler, which show substantial internal variation. These cases exemplify how different urban contexts contribute to the morphological diversity captured by the clustering results, without delving into historical or social specifics.
@@ 5.2.1 Marginal Effects of Urban Morphology on Modal Mobility
To show how different spatial configurations are associated with variations in the use of specific modes of transportation, we analyzed the relationship between urban morphology and modal usage through the estimated marginal effects. Figure 4 presents a heatmap displaying the normalized coefficients associated with each urban pattern, allowing for a visual interpretation of how certain morphological forms are linked to positive or negative deviations in the propensity for active, public, or private mobility compared to the overall average.
@@ 6 Discussion
This association is further supported by quantitative results across all cities via post hoc analysis (Table 1). For instance, comparing gridiron and organic patterns reveals statistically significant differences in private mobility (p-value = ), with a negative effect size () indicating a lower proportion of this mode in areas with a more structured street network. Similarly, the effects on active and public mobility are positive ( and , respectively), reinforcing the notion that spatial order in urban form tends to foster more sustainable transport behavior. These tendencies also appear in the comparison between cul-de-sac and gridiron patterns, highlighting the structural role of urban morphology in shaping residents’ modal decisions. Overall, the findings support the central hypothesis of the study, suggesting that urban form influences not only physical accessibility but also the functional orientation of mobility systems.
@@
Looking forward, future research should aim to deepen the theoretical and empirical understanding of how street network morphology shapes — and is shaped by — broader socio-spatial processes. While this study provided a scalable and transferable method to classify urban form and relate it to travel behavior, further work is needed to explore how these morphological patterns interact with topographic constraints, land use configurations, environmental risks, and infrastructural inequalities. In particular, we see potential in developing a critical and computational framework that bridges urban theory with topological analysis and open-source data tools. This would enable the construction of comparative morphologies across diverse geographies — especially in cities of the Global South — and support planning efforts focused on equitable, resilient, and sustainable urban transformations.

### v2

@@ Abstract
Urban morphology has long been recognized as a factor shaping human mobility, yet comparative and formal classifications of urban form across metropolitan areas remain limited. Building on theoretical principles of urban structure and advances in unsupervised learning, we systematically classified the built environment of nine U.S. metropolitan areas using structural indicators such as density, connectivity, and spatial configuration. The resulting morphological types were linked to mobility patterns through descriptive statistics, marginal effects estimation, and post hoc statistical testing. Here we show that distinct urban forms are systematically associated with different mobility behaviors, such as reticular morphologies being linked to significantly higher public transport use (marginal effect = 0.49) and reduced car dependence (–0.41), while organic forms are associated with increased car usage (0.44), and substantial declines in public transport (–0.47) and active mobility (–0.30). These effects are statistically robust (), highlighting that the spatial configuration of urban areas plays a fundamental role in shaping transportation choices. Our findings extend previous work by offering a reproducible framework for classifying urban form and demonstrate the added value of morphological analysis in comparative urban research. The dataset and code are openly available, allowing replication and adaptation to other geographical contexts. These results suggest that urban form should be treated as a key variable in mobility planning and provide empirical support for incorporating spatial typologies into sustainable urban policy design.
@@
The physical structure of cities exerts a determining influence on the daily lives of their inhabitants, shaping the way they move, access services, and participate in urban dynamics. In this context, urban morphology stands as a crucial factor in the organization of mobility, an aspect increasingly relevant in global debates on walkability, transit-oriented development, and the 15-minute city paradigm. This morphology not only defines the accessibility and connectivity of a territory but also conditions the modal share of trips and, consequently, equitable access to the city’s work, educational, and social opportunities (Geurs and Wee, 2004).
Historically, urban and transport policies have tended to address mobility from a predominantly functionalist perspective, focused on the operational efficiency of transport systems. However, this approach has often underestimated how urban design can either foster or restrict certain modes of travel. While previous research has demonstrated that some urban forms promote active travel and public transit, whereas others foster car dependency (Ewing and Cervero, 2010), much less attention has been paid to how localized street network subpatterns shape travel behavior. This represents a critical gap in understanding how urban form influences equity and sustainability in mobility. Recent works have increasingly employed deep learning and graph embeddings to characterize urban street morphology (e.g., Boeing (2024); Wu et al. (2024); Chen et al. (2021)), yet few have systematically linked these morphological clusters to behavioral outcomes such as travel mode choice. Our contribution bridges this methodological and empirical gap.
In this context, the present study examines how the morphological configuration of street networks influences urban mobility patterns. It analyzes the structural attributes of road systems as captured by topological and spatial indicators (Bamakan et al., 2019; Cardillo et al., 2006), and specifically investigates the explanatory power of subpatterns within larger networks. The central research question is: to what extent do variations in local street network subpatterns explain differences in travel mode choice, beyond socioeconomic and land-use factors? Through a comparative approach across U.S. cities, the study provides empirical evidence that can inform more integrated urban and transport planning policies aimed at promoting more equitable, sustainable, and efficient mobility systems.
@@ 2 Urban Morphology and Spatial Structure
In this study, we adopt a network-based understanding of urban morphology, where the structure and geometry of the street network are treated as a proxy for the underlying form of the city (Porta et al., 2006; Boeing, 2019). We acknowledge that this is only one of many valid ways to study urban form and do not claim conceptual primacy. Rather, we recognize that form is multidimensional, and that our approach captures one of its structural expressions—specifically, the topological and geometric configuration of urban streets as they relate to patterns of movement and accessibility. This computational morphology perspective aligns with the emerging field of urban systems science, where cities are analyzed as complex adaptive networks (Batty, 2012; Barthélemy, 2011)
@@ 2.1 Characterization of Urban Morphology: Fundamental Metrics and Dimensions
From this perspective, centrality metrics provide an additional layer of analysis that enables the examination of how accessibility and connectivity are distributed across different street fabrics. When interpreted through a functional lens, these measures help identify not only local connection patterns but also global articulation dynamics within the network. Degree centrality (Freeman, 1978), by estimating the number of direct connections each node has relative to the total, offers a first approximation of its level of integration; building on this, closeness centrality (Musiał et al., 2009; Wasserman and Faust, 1994) broadens the scope by considering the average distances between nodes, allowing for the identification of strategic points with greater potential access to the overall system. Complementarily, betweenness centrality as elaborated by Freeman (1978), introduces a critical dimension by revealing which nodes tend to lie along the shortest paths between other pairs, highlighting their role as articulators of general flow. A similar logic can be applied to the network’s segments through edge betweenness centrality (Lu and Zhang, 2013), which helps identify links with a high structural weight in overall connectivity. Although these metrics were not part of the empirical analysis developed in this study, they represent a complementary analytical dimension that can enrich the study of spatial configurations and the structural understanding of urban networks. These were excluded from the main clustering pipeline to prioritize variables with stable distributions across diverse urban contexts and to avoid multicollinearity issues in PCA.
@@ 3 Methodology for Theoretical Classification and Clustering of Urban Patterns
To conduct the morphological analysis of the street networks in the selected cities, the urban territory was divided into administrative units defined by the Census Bureau, known as census tracts (U.S. Census Bureau, 2024). This territorial segmentation enables a precise and consistent characterization of urban morphology, facilitating the integration of topological and spatial properties with complementary data such as mobility indicators available at the tract level. This approach enables systematic comparisons across urban areas while preserving the granularity needed to capture internal heterogeneity.
Figure 2 presents the overall methodological framework employed in this study. The analytical process begins with the extraction of spatial and topological metrics from street networks, which are subsequently used for theoretical classification through a Multi-Attribute Decision Making (MADM) approach. In parallel, unsupervised clustering techniques combining PCA and K-means are applied to reveal emergent subpatterns within the theoretically defined typologies, enabling a comprehensive characterization of urban morphology at multiple scales.
Following this framework, urban morphology is analyzed based on the variables previously defined and summarized in Table LABEL:tab:urban_metrics, which group key spatial and topological characteristics relevant to describing street networks. Building on this foundation, a systematic and replicable classification method is applied to quantify the degree of correspondence between each urban unit and the theoretical profiles of urban patterns established in the literature (Section 2.2). This method employs a Multi-Attribute Decision Making (MADM) (Triantaphyllou, 2000) framework that assigns weighted scores and penalties according to how closely the observed morphological attributes align with the characteristic values of each pattern. Within this framework, three types of intervals are identified. The optimal interval includes values that are highly characteristic and representative of a specific pattern, the moderate interval comprises values compatible with the pattern but less distinctive or potentially overlapping with others, and the critical or penalizing interval encompasses values that significantly contradict the defining features of the pattern. For example, a very high proportion of dead-end streets is considered critical when evaluating the Gridiron pattern, triggering a penalty within the scoring scheme. To quantify the correspondence between an urban area and each profile, a weighted scoring mechanism with penalties is applied. Let denote the value of dimension , and the weight assigned based on its relevance to the pattern. The partial score for each dimension is determined by the interval in which falls.
@@ 3.2 Cluster-Based Pattern and Subpattern Classification
Although Table LABEL:tab:urban_metrics defined conceptual groups to organize morphological properties, the PCA analysis shows that the empirical expression of these categories may vary depending on urban scale and internal heterogeneity. Nevertheless, the extracted dimensions tend to preserve the structural logic of the theoretical classification, suggesting a robust correspondence between conceptual patterns and actual urban configurations, without implying a rigid segmentation. Based on these dimensions, clustering techniques were applied to identify predominant morphological patterns, selecting the optimal number of clusters using metrics such as the silhouette score. Although the resulting groupings reflect configurations consistent with the theoretical framework, they should be understood as flexible structures in complex urban environments such as those with large territorial extensions or multilevel developments. Peripheral subgroups may emerge that, despite deviating from the cluster centroid, retain fundamental structural properties that justify their classification. All analyses were conducted in Python 3.11 using the OSMnx and scikit-learn libraries. The complete reproducibility pipeline is available at the project’s GitHub page.
4 Selection of Case Studies
The selection of cities was guided by the principle that each case should clearly represent one of the major morphological archetypes of urban street networks: grid, organic, hierarchical (cul-de-sac), and hybrid. This typological contrast enables both the validation of the proposed spatial and topological metrics across distinct structural regimes and the exploration of transitions between idealized and mixed forms. Each city was selected as a canonical or transitional representative of its morphological class, ensuring both historical depth and diversity of urban contexts.
4.1 Grid Pattern: Philadelphia and Salt Lake City
Philadelphia (Pennsylvania) represents one of the earliest and most influential grid plans in North America. Conceived by William Penn and Thomas Holme in 1682, its rectilinear layout with wide, uniform streets and five public squares reflects an intentional model for an ordered and fire-resistant city (Southworth and Ben-Joseph, 2003; Reps, 1965; Jackson, 1985). Conversely, Salt Lake City (Utah) illustrates a grid derived from ideological planning principles, following Joseph Smith’s Plat of Zion and implemented by Brigham Young in 1847 (Greenspan, 2016). The city’s grid, aligned with the cardinal axes and centered on the Temple, features unusually wide streets (approximately 40 m), designed to accommodate ox-drawn carts, thus reinforcing a symbolic and functional vision of order and expansiveness.
4.2 Organic and Early Hybrid Patterns: Boston and Santa Fe
Boston (Massachusetts) exemplifies an organically evolved urban morphology, shaped by topography and historical accretion rather than centralized planning (Warner Jr., 1978; Southworth and Owens, 1993). The street system of the Shawmut Peninsula follows preexisting paths and contours, producing a highly irregular network with variable block geometry and intersection angles. Although later interventions sought to regularize the city, most notably in the Back Bay expansion, (Whitehill and Kennedy, 2000) the influence on the historic core remains a paradigmatic organic pattern. Similarly, Santa Fe (New Mexico) combines planned and adaptive principles: founded under the Spanish Laws of the Indies with a central plaza (Crouch et al., 1982), its subsequent growth accommodated the terrain and existing trails, yielding a semi-regular yet curvilinear form typical of early hybrid developments.
4.3 Hierarchical Pattern (Cul-de-Sac): Peachtree City, Cary, and Chandler
This group encompasses suburban cities whose layouts are dominated by dendritic hierarchies and dead-end streets. Peachtree City (Georgia), established in 1959, is a model of mid-century planned suburbanism structured into self-contained villages (Garvin, 2002). Its defining feature is an extensive system of cul-de-sacs and loops connected by over 160 km of light-vehicle paths (O’Toole, 2009). Cary (North Carolina) similarly reflects the proliferation of cul-de-sacs through Planned Unit Developments (PUDs); however, recent urban plans encourage higher street connectivity (Town of Cary, 2020). Chandler (Arizona), while originally agricultural and gridded, experienced massive postwar suburban expansion that replaced large blocks with residential subdivisions characterized by dendritic and cul-de-sac structures (Southworth and Ben-Joseph, 2003).
4.4 Hybrid and Overlapping Patterns: Charleston and Fort Collins
Charleston (South Carolina) and Fort Collins (Colorado) illustrate transitional and composite urban fabrics where multiple planning logics coexist. Charleston originated from the Grand Model of 1672, introducing a colonial grid that was later modified by coastal topography and incremental development (Reps, 1965). The resulting fabric juxtaposes rectilinear order with organic irregularity. Fort Collins, in turn, presents a postwar superposition of morphological layers: a compact, walkable grid in the historic core, surrounded by curvilinear suburban subdivisions of the automobile era (Handy et al., 2003).
In summary, the nine selected cities constitute a coherent comparative framework to test and validate the proposed morphological metrics. Together, they span the full spectrum of urban form, from regular, high-connectivity grids to dendritic, low-connectivity networks while also including transitional configurations that reveal the fluid boundaries between planning paradigms. This diversity allows assessing the model’s robustness across contrasting structural conditions and exploring how different spatial organizations relate to connectivity, entropy, and modal mobility patterns at the urban scale.
5 Classification and Treatment of Urban Mobility
In order to analyze the relationship between urban morphology and mobility patterns, data on street space usage are collected and classified based on information provided by the U.S. Census Bureau (2020). Specifically, we used ACS 5-year estimates (2017–2021) of commuting mode share at the census tract level, normalized by working-age population. This information is organized following the ABC of Mobility framework Prieto-Curiel and Ospina (2024), which classifies travel modes into three categories based on their function and means of transport. Active mobility includes non-motorized forms such as walking and cycling, public mobility comprises trips made using shared or collective transport systems, and private mobility refers to the use of individual motorized vehicles. This classification serves to structure the analysis of modal distribution in a clear and comparable way. Based on this classification, mobility patterns are examined from a dual perspective. On the one hand, the disaggregated approach makes it possible to assess the specific contribution of each transport mode and its potential correlation with particular morphological attributes. On the other hand, the aggregated analysis provides a general overview of the modal composition of each urban unit, allowing for systematic comparisons across different spatial contexts. This methodological approach not only enables the characterization of transport mode distribution but also facilitates the exploration of potential associations between the built environment and everyday mobility practices.
@@ 6.1 Morphological Pattern Characterization in Selected Cities
Figure 3 presents selected examples of morphological categories and sub-patterns in cities like Salt Lake City, Boston, and Chandler, which show substantial internal variation. These cases exemplify how different urban contexts contribute to the morphological diversity captured by the clustering results, without delving into historical or social specifics.
@@ 6.2.1 Marginal Effects of Urban Morphology on Modal Mobility
To show how different spatial configurations are associated with variations in the use of specific modes of transportation, we analyzed the relationship between urban morphology and modal usage through estimated marginal effects. These effects quantify the deviation of each urban pattern from the overall modal share baseline, calculated as the difference between the mean modal usage within each typology and the global average across all urban units. To facilitate comparison across mobility modes with different scales, the marginal effects are normalized by the standard deviation of each variable, yielding standardized effect sizes that indicate the magnitude and direction of the association. Figure 5 presents a heatmap displaying these normalized coefficients for each urban pattern, allowing for a visual interpretation of how certain morphological forms are linked to positive or negative deviations in the propensity for active, public, or private mobility.
@@ 7 Discussion
This association is further supported by quantitative results across all cities via post hoc analysis (Table 1). For instance, comparing gridiron and organic patterns reveals statistically significant differences in private mobility (p-value = ), with a negative effect size () indicating a lower proportion of this mode in areas with a more structured street network. Similarly, the effects on active and public mobility are positive ( and , respectively), reinforcing the notion that spatial order in urban form tends to foster more sustainable transport behavior. These tendencies also appear in the comparison between cul-de-sac and gridiron patterns, highlighting the structural role of urban morphology in shaping residents’ modal decisions. Overall, the findings support the central hypothesis of the study, suggesting that urban form influences not only physical accessibility but also the functional orientation of mobility systems. While this study focuses on U.S. cities, future research should expand to Global South contexts, where morphological diversity and informal development patterns may yield distinct mobility relationships.
@@
Looking forward, future research should aim to deepen the theoretical and empirical understanding of how street network morphology shapes — and is shaped by — broader socio-spatial processes. While this study provided a scalable and transferable method to classify urban form and relate it to travel behavior, further work is needed to explore how these morphological patterns interact with topographic constraints, land use configurations, environmental risks, and infrastructural inequalities. In particular, we see potential in developing a critical and computational framework that bridges urban theory with topological analysis and open-source data tools. This would enable the construction of comparative morphologies across diverse geographies — especially in cities of the Global South — and support planning efforts focused on equitable, resilient, and sustainable urban transformations. By coupling computational morphology with behavioral data, this work contributes to the emerging paradigm of morphological informatics, advancing the use of data-driven typologies to inform urban mobility policy.
8 Conclusions
The quantitative analysis of street networks across the selected cities confirms that urban morphology constitutes a structural determinant in shaping mobility patterns. The unsupervised classification methodology not only reproduced the canonical typologies of grid, organic, and hierarchical (cul-de-sac) networks, along with their hybrid variants, but also provided a significant contribution by revealing internal variability within each category. This finer resolution demonstrates that theoretical typologies are not monolithic constructs, but rather continua that encompass measurable differences in density, connectivity, and geometric configuration, fulfilling the objective of quantifying the road network structure in detail.
A systematic and statistically significant correspondence was observed between urban form and modal split. Grid-based structures, characterized by high connectivity and permeability, consistently support higher shares of active and public transport modes. In contrast, dendritic cul-de-sac systems and irregular organic layouts are associated with a substantially greater dependence on private vehicles. The magnitude of these associations is underscored by the post hoc analyses, where the differences in automobile use between grid and organic morphologies exhibit large effect sizes, emphasizing their practical relevance. Taken together, the results reveal consistent trends linking network connectivity to transport sustainability, thereby addressing the core research objectives. Urban form thus emerges not as a passive spatial backdrop, but as an active mechanism that enables or constrains specific travel behaviors.
Overall, the findings substantiate that the configuration of the street network exerts a significant influence on urban mobility, supporting the central propositions of urban form theory. Nevertheless, the strength of these effects varies across contexts, indicating that additional factors—such as socioeconomic conditions, topography, and planning policies—mediate the relationship between morphology and mobility. This heterogeneity highlights the need for future multiscalar approaches that integrate morphological structure with broader territorial and social dynamics. Beyond validating the structural importance of urban form, this study opens avenues for comparative research in more complex and unequal urban environments, particularly in Latin American cities, where spatial fragmentation and mobility inequities pose critical challenges for sustainable urban transitions.